# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xxxx

### v2

@@ Abstract
Fourier Analysis Network (FAN) was recently proposed as a simple way to improve neural network performance by replacing part of Rectified Linear Unit (ReLU) activations with sine and cosine functions. Although several studies have reported small but consistent gains across tasks, the underlying mechanism behind these improvements has remained unclear. In this work, we show that only the sine activation contributes positively to performance, whereas the cosine activation tends to be detrimental. Our analysis reveals that the improvement is not a consequence of the sine function’s periodic nature; instead, it stems from the function’s local behavior near $x=0$, where its non-zero derivative mitigates the vanishing-gradient problem. We further show that FAN primarily alleviates the dying-ReLU problem, in which a neuron consistently receives negative inputs, produces zero gradients, and stops learning. Although modern ReLU-like activations, such as Leaky ReLU, GELU, and Swish, reduce ReLU’s zero-gradient region, they still contain input domains where gradients remain significantly diminished, contributing to slower optimization and hindering rapid convergence. FAN addresses this limitation by introducing a more stable gradient pathway. This analysis shifts the understanding of FAN’s benefits from a spectral interpretation to a concrete analysis of training dynamics, leading to the development of the Dual-Activation Layer (DAL), a more efficient convergence accelerator. We evaluate DAL on three tasks: classification of noisy sinusoidal signals versus pure noise, MNIST digit classification, and Electrocardiogram (ECG)-based biometric recognition. In all cases, DAL models converge faster and achieve equal or higher validation accuracy compared to models with conventional activations.
@@
Convolutional neural network, deep learning, Fourier Analysis Network, machine learning, activation functions.
@@ I Introduction
In this paper, we clarify the mechanism underlying FAN through simple 1D signal-classification experiments that distinguish between (a) noisy sinusoids and (b) pure noise. Our main findings are:
@@
    Only the sine activation contributes positively. We show that the sine activation consistently improves network performance, whereas the cosine activation tends to degrade it. We further demonstrate that the improvement originates from the sine function’s non-zero derivative at the origin rather than its periodic nature, providing a critical pathway for continuous optimization. Building on this insight, we propose the Dual Activation Layer (DAL), denoted as (f,g,a:b)(f,g,a{:}b), which mixes two activation functions $f$ and $g$ in proportion a:ba{:}b. Here, $f$ is ReLU (or a similar variant such as Leaky ReLU, GELU, or Swish), and $g$ is sine (or another function with a similar shape near $x=0$, such as tanhtanh).
@@
Finally, we apply DAL to real-world tasks — handwritten digit classification (MNIST) and electrocardiogram-based biometric recognition (ECG-ID) — and observe consistent acceleration in convergence across both problems. In the ECG-ID task, we observed not only faster convergence but also an improvement in the network’s final performance, perhaps because it is a more complex problem.
@@ II Related works
Other studies have explored variants and extensions of FAN within transformer architectures and KAN models [Tkhai2025, Eslamian2025, Wang2025, Tan2025]. Wang et al. [Wang2025] report that incorporating FAN into a transformer increases the mean accuracy for EEG-based emotion recognition from 67.78% to 70.30%. However, the large and highly overlapping standard deviations (9.69 vs. 9.89) raise doubts about the statistical significance of this improvement. Similarly, Tan et al. [Tan2025] show that a Fourier-based KAN improves the average DICE score on a hepatic vessel segmentation task from 66.57% to 67.57%, yet no statistical analysis is provided to assess whether this difference is meaningful. This uncertainty is further reinforced by the findings of Eslamian et al. [Eslamian2025], who evaluated multiple KAN variants across diverse tasks and concluded that FAN-based KANs do not consistently achieve state-of-the-art performance. Their results suggest that the gains reported in some studies may fall within natural performance variability rather than reflecting a systematic or robust advantage.
@@ III-A Problem Setup
Figure 1 illustrates two examples of each signal type (nv01.py). In our experiments, we compare two network architectures, CNNA and CNNB, on the task of classifying each signal as sine or noise. For each trial, we generate 100 signals from each class, shuffle them, and split the dataset in half: 50 for training and 50 for testing. Both networks are trained and evaluated on exactly the same set of signals. Training lasts for $M$ epochs, and the entire procedure is repeated $N$ times, yielding $N$ paired test accuracies. We then apply paired hypothesis tests to assess whether the differences between CNNA and CNNB are statistically significant.
@@ III-C Reproducibility
All programs associated with this article are available at a shared folder  [^1] https://drive.google.com/drive/folders/1ScArmAPIaOvxzbrEeIi-4HFmWnjstaNY, and their filenames are referenced throughout the text. Portions of the code and text were developed with the assistance of artificial intelligence tools.
@@ IV Experiments and Results
IV-A (ReLU, sin, cos, 6:1:1) vs. ReLU-only
@@
We begin our experiments by evaluating the original idea behind FAN: combining a conventional activation function (such as ReLU) with sine and cosine activations. To this end, we compare a network architecture that uses ReLU, sine, and cosine in a 6:1:1 ratio against a model that uses ReLU exclusively. We adopted the 6:1:1 proportion — rather than the originally suggested 2:1:1 — because it yielded slightly better results in our preliminary tests. We performed N=50N{=}50 independent training runs, each with M=10M{=}10 epochs, obtaining the following results (nv17.py):
Mean Accuracy (CNNA  ReLU/sin/cos 6:1:1): 0.8340
Mean Accuracy (CNNB  ReLU): 0.7826
P-value (paired test): 0.0026
@@
These results indicate, with strong statistical significance $(p{=}0.26\%)$, that replacing a fraction of the ReLU activations with sine and cosine leads to improved performance. However, a key question remains: Is the improvement driven primarily by the sine activation, by the cosine activation, or by the combination of both? The experiments in the following sections address this question directly.
IV-B (ReLU, cos, 6:2) vs. ReLU-only
@@
We now evaluate the effect of using only the cosine activation, completely removing the sine activation. To this end, we compare (ReLU, cos, 6:2) with the baseline ReLU-only. Running N=50N{=}50 trials, each with M=10M{=}10 training epochs, we obtained (nv19.py):
Mean Accuracy (CNNA  ReLU/cos 6:2): 0.7608
Mean Accuracy (CNNB  ReLU): 0.7906
In contrast to the performance gains observed when sine activations are included, replacing a portion of the ReLU activations with cosine leads to a degradation in performance. This result indicates that the cosine activation does not contribute positively — indeed, it tends to harm network performance.
IV-C (ReLU, sin, 6:2) vs. ReLU-only
@@
Next, we evaluate the effect of combining ReLU with the sine activation, this time excluding cosine entirely. We compare (ReLU, sin, 6:2) against the baseline ReLU-only. Running N=100N{=}100 trials, each with M=10M{=}10 training epochs, we obtain (nv21.py):
Mean Accuracy (CNNA  ReLU/sin 6:2): 0.8153
Mean Accuracy (CNNB  ReLU): 0.7787
P-value for the accuracy difference: 0.0035
@@
This statistically significant improvement ($p{=}0.35\%$) shows that the sine activation is the primary contributor to the performance gains previously attributed to FAN. Figure 2 illustrates the average evolution of the performance metrics over the 10 training epochs, aggregated across all 100 runs. The model using (ReLU, sin, 6:2) converges more quickly and consistently outperforms the pure-ReLU model throughout training.
@@
Figure 2: Evolution of performance metrics (averaged over N=100N{=}100 runs) for (ReLU, sin, 6:2) and ReLU-only. The former exhibits consistently superior performance.
IV-D (ReLU, sin, 6:2) vs. ReLU-only with $M=60$ Epochs
@@
The results from the previous section suggested that (ReLU, sin, 6:2) achieved higher test accuracy than pure-ReLU. However, this advantage disappears when the training is extended. Repeating the same experiment N=100N{=}100 times, now with M=60M{=}60 epochs, we obtain (Figure 3, nv21f.py):
Mean Accuracy (CNNA  ReLU/sin 6:2): 0.8579
Mean Accuracy (CNNB  ReLU): 0.8572
P-value for the accuracy difference: 0.9127
With a sufficiently large number of epochs, the two models reach essentially identical final accuracy. Nonetheless, as shown in Figure 3, the model using (ReLU, sin, 6:2) converges noticeably faster, achieving high accuracy earlier than the pure-ReLU model. In many problems, therefore, the ReLU–sine combination tends to speed up training even when the eventual final accuracy remains unchanged. Such an acceleration is valuable in practice and should not be overlooked.
The 6:2 ratio was not optimized, and different proportions may yield better results. When we examine the test losses, both models reach a minimum and then begin to worsen, most likely due to overfitting. We also note that, in some more complex tasks and architectures, the (ReLU, sin) combination does achieve higher final accuracy than plain ReLU, even after extensive training (see Subsection V-B)
@@
Figure 3: Evolution of performance metrics averaged over N=100N{=}100 runs across M=60M{=}60 epochs for (ReLU, sin, 6:2) and ReLU-only.
IV-E (ReLU, sin, cos, 6:1:1) vs. (ReLU, sin, 6:2)
To assess whether the proposed DAL architecture (ReLU, sin, 6:2) indeed outperforms the original FAN architecture (ReLU, sin, cos, 6:1:1) we directly compared the two models. We conducted N=50N{=}50 trials, each with M=10M{=}10 training epochs, and evaluated the evolution of the performance metrics throughout training (Figure 4, nva20.py).
We observed that the DAL (ReLU, sin, 6:2) model converges faster than the model based on the original FAN scheme (ReLU, sin, cos, 6:1:1). This result reinforces our earlier conclusion: incorporating the cosine activation does not improve network performance and, in practice, tends to hinder convergence. By demonstrating that the cosine function is superfluous, the DAL is proposed as the necessary and sufficient layer structure to achieve the rapid convergence observed in FANs.
Figure 4: Evolution of performance metrics for the (ReLU, sin, cos, 6:1:1) and (ReLU, sin, 6:2) architectures. The latter converges more rapidly than the former.
IV-F Non-periodic Activations
@@
We evaluated (ReLU, $f$, 6:2) for $f{\in}$ {TSine, tanh, linear}, and compared the results with (ReLU, sine, 6:2). Running N=50N{=}50 tests with M=10M{=}10 epochs each, we obtained the results shown in Figures 6, 7, and 8 (nv23.py, nvb01.py, nvb02.py). Networks using TSine and tanh achieved performance levels virtually identical to those obtained with sine, whereas the linear activation exhibited slightly slower convergence.
Figure 6: (ReLU, TSine, 6:2) vs. (ReLU, sine, 6:2).
Figure 7: (ReLU, tanh, 6:2) vs. (ReLU, sine, 6:2)
Figure 8: (ReLU, linear, 6:2) vs. (ReLU, sine, 6:2)
Since the improvements persist even when sine is replaced by non-periodic functions that nonetheless resemble sine within the approximate interval $x{\in}[-1.5,+1.5]$, we conclude that the benefit does not arise from periodicity. Instead, it comes from the local shape of the function around x=0x{=}0, which helps accelerate training.
By computing the mean and standard deviation of the values entering the activation functions across different network layers, we obtained $\mu{\approx}0$ and $\sigma{\in}[0.2,0.9]$ (nv90.py). Thus, assuming the values follow a normal distribution, most inputs to the activation functions fall within the interval $x{\in}[-1.5,+1.5]$.
IV-G Dead Neurons
@@
To test this hypothesis, we trained the models (ReLU, sine, 6:2) and ReLU-only, each N=20N{=}20 times for M=20M{=}20 epochs (nv81.py). We then computed and printed, for each trained network, the average percentage of dead neurons per layer across the 20 runs. Table I shows that (ReLU, sine, 6:2) exhibits substantially fewer dead neurons than ReLU alone. The alternative activation (sine) provides an additional gradient pathway, reducing the number of dead neurons. Incorporating sine yields a more robust gradient flow and lowers the likelihood of neuron death during training. A direct implication of this reasoning is that TSine should be avoided in practice: it contains wide regions with zero gradient, which can also lead to dead neurons, much like ReLU.
Layer (ReLU, sine, 6:2) Only ReLU
@@
TABLE I: Percentage of dead neurons per layer for the models (ReLU, sine, 6:2) and only ReLU.
IV-H Modern Activation Functions
@@
Figure 11: (Leaky Relu, sin, 7:1) converges faster than pure Leaky Relu.
@@
V Real-World Applications
V-A MNIST Classification
To assess whether the conclusions drawn from the synthetic 1D classification problem generalize to a more realistic task, we evaluated handwritten digit classification on the MNIST dataset using convolutional neural networks with activation (ReLU, sine, 6:2) and pure ReLU. Both models share the following architecture:

### v3

@@ Abstract
Fourier Analysis Network (FAN) was recently proposed as a simple way to improve neural network performance by replacing part of Rectified Linear Unit (ReLU) activations with sine and cosine functions. Although several studies have reported small but consistent gains across tasks, the underlying mechanism behind these improvements has remained unclear. In this work, we show that only sine contributes positively to performance, whereas cosine tends to be detrimental. Our analysis reveals that the improvement is not a consequence of the sine function’s periodic nature; instead, it stems from the function’s local shape in the neighborhood of x=0x{=}0. We further show that sine activation primarily alleviates the dying-ReLU problem, in which a neuron consistently receives negative inputs, produces zero gradients, and stops learning. Although modern ReLU-like activations eliminate zero-gradient region, they still contain input domains where gradients remain diminished, hindering rapid convergence. Sine activation addresses this limitation by introducing a more stable gradient pathway. This analysis leads to the development of the Dual-Activation Layer (DAL), a more efficient convergence accelerator. We evaluate DAL on three tasks: classification of noisy sinusoidal signals versus pure noise, MNIST digit classification, and Electrocardiogram (ECG)-based biometric recognition. In all instances, DAL-based models exhibit faster convergence rates relative to models using conventional activations. Furthermore, DAL achieved slightly higher final test accuracy across all three tasks. This improvement was statistically significant for MNIST and ECG-ID; however, while a performance gain was observed in the simpler sinusoidal classification task, it did not reach the threshold for statistical significance.
@@
Fourier Analysis Network, activation functions, dying-ReLU problem, Dual-Activation Layer, neural network training dynamics, convolutional neural network.
@@ I Introduction
In this paper, we clarify the mechanism underlying FAN through simple 1D signal-classification experiments that distinguish between noisy sinusoids and pure noise. Our main findings are:
@@
    Only the sine activation contributes positively. We show that the sine activation consistently improves network performance, whereas the cosine activation tends to degrade it. We further demonstrate that the improvement originates from the sine function’s non-zero derivative at the origin rather than its periodic nature, providing a critical pathway for continuous optimization. Building on this insight, we propose the Dual-Activation Layer (DAL), denoted as (f,g,a:b)(f,g,a{:}b), which mixes two activation functions $f$ and $g$ in proportion a:ba{:}b. Here, $f$ is ReLU (or a similar variant), and $g$ is sine (or another function with a similar shape near $x=0$, such as tanhtanh). Figure 1 provides a simplified comparison between the ReLU-only, FAN, and DAL activation layers for a single-channel configuration. In multi-channel layers, this structure is replicated on a per-channel basis. Each activation function is statically assigned to a specific neuron throughout the model’s lifecycle, from training through to inference.
@@
Finally, we apply DAL to real-world tasks — handwritten digit classification (MNIST) and electrocardiogram-based biometric recognition (ECG-ID) — and observe consistent acceleration in convergence and statistically significant higher final accuracy across both problems.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 1: Schematic representation of activation layers composed of four neurons: (a) ReLU-only; (b) FAN (ReLU, sin, cos, 2:1:1); (c) DAL (ReLU, sin, 1:1).
@@ II Related works
Other studies have explored variants and extensions of FAN within transformer architectures and Kolmogorov-Arnold Networks (KAN) [Tkhai2025, Eslamian2025, Wang2025, Tan2025]. Wang et al. [Wang2025] report that incorporating FAN into a transformer increases the mean accuracy for electroencephalogram (EEG)-based emotion recognition from 67.78% to 70.30%. However, the large and highly overlapping standard deviations (9.69 vs. 9.89) raise doubts about the statistical significance of this improvement. Similarly, Tan et al. [Tan2025] show that a Fourier-based KAN improves the average Dice similarity coefficient on a hepatic vessel segmentation task from 66.57% to 67.57%, yet no statistical analysis is provided to assess whether this difference is meaningful. This uncertainty is further reinforced by the findings of Eslamian et al. [Eslamian2025], who evaluated multiple KAN variants across diverse tasks and concluded that FAN-based KANs do not consistently achieve state-of-the-art performance. Their results suggest that the gains reported in some studies may fall within natural performance variability rather than reflecting a systematic or robust advantage.
@@ III-A Problem Setup
Figure 2 illustrates two examples of each signal type (nv01.py). In our experiments, we compare two network architectures, CNNA and CNNB, on the task of classifying each signal as sine or noise. For each trial, we generate 100 signals from each class, shuffle them, and split the dataset in half: 50 for training and 50 for testing. Both networks are trained and evaluated on exactly the same set of signals. Training lasts for $M$ epochs, and the entire procedure is repeated $N$ times, yielding $N$ paired test accuracies. We then apply one-sided paired t-test 111scipy.stats.ttest_rel to assess whether the differences between CNNA and CNNB are statistically significant.
@@ III-C Reproducibility
All programs associated with this article are available at a shared folder  [^2] https://drive.google.com/drive/folders/1ScArmAPIaOvxzbrEeIi-4HFmWnjstaNY, and their filenames are referenced throughout the text.
@@ IV Experiments and Results
IV-A FAN(ReLU, sin, cos, 6:1:1) vs. ReLU-only, M=6M{=}6
@@
We begin our experiments by evaluating the original idea behind FAN: combining a conventional activation function (such as ReLU) with sine and cosine activations. To this end, we compare a FAN architecture that uses ReLU, sine, and cosine in a 6:1:1 ratio against a model that uses ReLU exclusively. We adopted the 6:1:1 proportion — rather than the originally suggested 2:1:1 — because it yielded slightly better results in our preliminary tests. We performed N=100N{=}100 independent training runs, each with M=6M{=}6 epochs, obtaining the following results (nv17.py):
Mean Test Accuracy (CNNA - relu/sin/cos 6:1:1): 0.7638
Mean Test Accuracy (CNNB - relu): 0.7378
One-sided (>) paired p-value: 0.0150
@@
These results indicate, with statistical significance (p=0.0150)(p{=}0.0150), that replacing a fraction of the ReLU activations with sine and cosine leads to improved performance. However, a key question remains: Is the improvement driven primarily by the sine activation, by the cosine activation, or by the combination of both? The experiments in the following sections address this question directly.
IV-B (ReLU, cos, 6:2) vs. ReLU-only, M=6M{=}6
@@
We now evaluate the effect of using only the cosine activation, completely removing the sine activation. To this end, we compare (ReLU, cos, 6:2) with the baseline ReLU-only. Running N=50N{=}50 trials, each with M=6M{=}6 training epochs, we obtained (nv19.py):
Mean Test Accuracy (CNNA - relu/cos 6:2): 0.6804
Mean Test Accuracy (CNNB - relu): 0.7600
One-sided (<) paired p-value: 0.0000
Substituting a portion of ReLU activations with cosine functions leads to a significant degradation in performance (p<0.001p{<}0.001). This result suggests that the cosine activation negatively impacts network efficacy in this configuration.
IV-C DAL(ReLU, sin, 6:2) vs. ReLU-only, M=6M{=}6
@@
Next, we evaluate the effect of combining ReLU with the sine activation, this time excluding cosine entirely. We compare DAL (ReLU, sin, 6:2) against the baseline ReLU-only. Running N=100N{=}100 trials, each with M=6M{=}6 training epochs, we obtain (nv21.py):
Mean Test Accuracy (CNNA - relu/sin 6:2): 0.8090
Mean Test Accuracy (CNNB - relu): 0.7338
One-sided (>) paired p-value: 0.0000
@@
This highly significant improvement (p<0.001p{<}0.001) shows that the sine activation is the primary contributor to the performance gains previously attributed to FAN. Figure 3 illustrates the average evolution of the performance metrics over the 6 training epochs, aggregated across all 100 runs. The model using DAL converges more quickly and consistently outperforms the pure-ReLU model throughout training.
@@
Figure 3: Evolution of performance metrics (averaged over N=100N{=}100 runs) for DAL(ReLU, sin, 6:2) and ReLU-only. The former exhibits consistently superior performance.
While Keras and TensorFlow utilize Glorot (Xavier) Uniform as the default initialization, we extended our experiments to include He (Kaiming) Normal and Uniform distributions. In all cases, the DAL architecture demonstrated superior convergence rates compared to the ReLU-only baseline, regardless of the initialization policy employed.
IV-D DAL(ReLU, sin, 6:2) vs. ReLU-only, M=60M{=}60
@@
The advantage of DAL over pure-ReLU diminishes when the training is extended. Repeating the same experiment N=100N{=}100 times, now with M=60M{=}60 epochs, we obtain (Figure 4, nv21f.py):
Mean Test Accuracy (CNNA - relu/sin 6:2): 0.8642
Mean Test Accuracy (CNNB - relu): 0.8583
One-sided (>) paired p-value: 0.1137
The final accuracy achieved by the DAL model is higher than that of the ReLU-only baseline. However, the resulting $p$-value of 0.1137 is insufficient to conclude that the DAL significantly outperforms the baseline in terms of final performance. On the other hand, the DAL architecture exhibits a significantly higher convergence rate, as illustrated in Figure 4. Such an acceleration is of considerable practical value and should not be overlooked. The 6:2 ratio was not optimized, and different proportions may yield better results.
@@
Figure 4: Evolution of performance metrics averaged over N=100N{=}100 runs across M=60M{=}60 epochs for DAL(ReLU, sin, 6:2) and ReLU-only.
V Discussion
V-A Non-periodic Activations
@@
We evaluated (ReLU, $f$, 6:2) for $f{\in}$ {TSine, tanh, linear}, and compared the results with (ReLU, sin, 6:2). Running N=50N{=}50 tests with M=10M{=}10 epochs each, we obtained the results shown in Figures 6, 7, and 8 (nv23.py, nvb01.py, nvb02.py). Networks using TSine and tanh achieved performance levels virtually identical to those obtained with sine, whereas the linear activation exhibited slightly slower convergence.
Figure 6: Comparison of (ReLU, TSine, 6:2) vs. (ReLU, sin, 6:2) activation configurations.
Figure 7: Comparison of (ReLU, tanh, 6:2) vs. (ReLU, sin, 6:2).
Figure 8: (ReLU, linear, 6:2) vs. (ReLU, sin, 6:2)
Since the improvements persist even when sine is replaced by non-periodic functions that nonetheless resemble sine within the approximate interval $x{\in}[-1.5,+1.5]$, we conclude that the benefit does not arise from periodicity. Instead, it comes from the local shape of the function around x=0x{=}0.
Most neural network frameworks are designed to perform optimally when the input data are normalized. Consequently, most practical implementations include a pre-processing step that normalizes the inputs. In addition, popular weight initialization methods are designed to preserve an approximately standard normal distribution across the internal layers. Under this assumption, most values entering the activation functions are expected to lie roughly within the interval $x\in[-1.5,+1.5]$, which highlights the importance of the activation function’s shape within this range. By empirically computing the statistics of the pre-activation values in the program used in this section, we obtained $\mu\approx 0$ for all layers, $\sigma\approx 0.45$ for the convolutional layers, and $\sigma\approx 0.85$ for the dense layer (nv90.py). These results indicate that, in practice, the standard deviation is smaller than one, further reinforcing the importance of the activation function’s behavior in the interval $x\in[-1.5,+1.5]$.
V-B Dead Neurons
@@
To test this hypothesis, we trained the models DAL (ReLU, sin, 6:2) and ReLU-only, each N=20N{=}20 times for M=20M{=}20 epochs (nv81.py). We then computed and printed, for each trained network, the average percentage of dead neurons per layer across the 20 runs.
Table I shows that DAL (ReLU, sin, 6:2) exhibits substantially fewer dead neurons than ReLU alone. The alternative activation (sine) provides an additional gradient pathway, reducing the number of dead neurons. Incorporating sine yields a more robust gradient flow and lowers the likelihood of neuron death during training. Consequently, given two architectures with identical neuron counts, the DAL model maintains a higher population of active neurons in practice than the ReLU-only model. Utilizing DAL is therefore functionally equivalent to increasing model capacity without expanding the actual parameter count.
An implication of this reasoning is that TSine should be avoided in practice: it contains wide regions with zero gradient, which can also lead to dead neurons, much like ReLU.
Layer (ReLU, sin, 6:2) Only ReLU
@@
TABLE I: Percentage of dead neurons per layer for the models (ReLU, sin, 6:2) and only ReLU.
V-C Modern Activation Functions
@@
Figure 11: (Leaky Relu, sin, 7:1) converges faster than Leaky Relu-only.
@@
VI Real-World Applications
VI-A MNIST Classification
To assess whether the conclusions drawn from the synthetic 1D classification problem generalize to a more realistic task, we evaluated handwritten digit classification on the MNIST dataset using convolutional neural networks with activation DAL (ReLU, sin, 6:2) and pure ReLU. Both models share the following architecture: